OpenStudy (anonymous):

Does anybody know how to convert this into a polar graph? x+5y=8

OpenStudy (anonymous):

I've genuinely tried everything and still couldn't get an answer.

OpenStudy (john_es):

Try this, \[x=r\cos\theta\\ y=r\sin\theta\]

OpenStudy (john_es):

And then solve for r.

OpenStudy (anonymous):

\[r \cos \theta + 8 + r \sin \theta - 5\sin \theta = 8\], I assume?

OpenStudy (anonymous):

rsintheta/5sintheta=8*